OpenStudy (anonymous):

0.7x+0.4(20)=0.5(x+20)

OpenStudy (anonymous):

10

OpenStudy (anonymous):

do you just distribute

OpenStudy (anonymous):

0.7x+0.4(20)=0.5(x+20) 0.7x+0.4 * 20=0.5x+0.5 *20 0.7x-0.5x=0.5*20-0.4*20 0.2x=10-8 0.2x=2 x=2/0.2 x=10

OpenStudy (anonymous):

Does it help?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

WELCOME